You bought a February Eurodollar futures at 98.30. A month later, you closed out the contract by selling the futures are 98.10. What is your profit/loss? [notional principal = $1,000,000 / contract maturity = 3 months]

Group of answer choices

1.70%

-$500

1.90%

+500
Answer #1

Ans : + 500

Euro dollar futures is used to hedge against adverse interest rate move

When one sold eurodollar futures it means he is locking interest rate. At 98.30 . implied interest rate is 100-98.3 = 1.7%

now at expiry eurodollar futures was trading at 98.1 , means implied rate = 100-98.1 =  1.9%

Thus one is profited by (1.9%-1.7%) = 0.2% i.e 20 basis points

Thus profit = 20*25$ = 500$